# ChartGalaxy: A Dataset for Infographic Chart Understanding and Generation

### v3

@@ 3.5 Data Statistics
ChartGalaxy contains 1,151,087 programmatically generated infographic charts and 104,519 real ones, covering 75 chart types, 330 chart variations, and 68 layout templates. Among the 75 chart types, the most frequently occurring ones are horizontal bar charts (11.7%), vertical bar charts (4.9%), and scatterplots (3.5%), reflecting their common usage in infographic-style presentations. Each chart type has up to 26 variations, capturing a wide range of styles in icon placement, color encoding, and data-label associations. We also observe variation in the frequency of layout template usage. For example, the template in Fig. 3(f) is the most frequently used, accounting for 6.9% of the synthetic charts. These differences are influenced by the structural flexibility of each template, their compatibility with various chart types, and their prevalence in real-world use cases. For more information on chart types, chart variations, and templates, please refer to Supp. B and Supp. C. Each infographic chart in the dataset is associated with tabular data, providing rich supervision for training and evaluation tasks.
@@ 4.1 Instruction Dataset for Infographic Chart Understanding
In this experiment, we construct an instruction dataset with ChartGalaxy to enhance model capabilities on infographic chart understanding. We validate its usefulness by fine-tuning two open-source LVLMs, demonstrating clear improvements on both public benchmarks and our independent evaluation set.
Dataset construction   To improve LVLMs’ data comprehension and visual understanding of infographic charts, we construct an instruction dataset comprising 443,455 question-answer pairs based on 70,248 charts randomly sampled from ChartGalaxy. The questions are classified into three types: 1) Text-based reasoning. We incorporate well-established question types from prior work, including open-form questions [41] and template-based questions [39]. These questions cover data identification (DI), data comparison (DC), data extraction with condition (DEC), and fact checking (FC). 2) Visual-element-based reasoning. We extend beyond purely text-based reasoning questions by incorporating visual elements from charts, such as icons (e.g., “What was the wine export value of [Uncaptioned image] in 2024?”). These questions require models to associate visual elements with their corresponding data values, thus testing their ability to conduct more complex cross-modal reasoning [67]. 3) Visual understanding. This type includes style detection (SD), visual encoding analysis (VEA, e.g., “What data dimension is encoded using different colors in this infographic chart?”), and chart classification (CC). The three types of questions evaluate a model’s ability to interpret visual design elements and underlying structural representations of the data. Detailed prompts and methods for generating question-answer pairs are provided in Supp. E.1. We also construct an independent, human-verified evaluation set containing 2,176 charts with 4,975 question-answer pairs for systematic evaluation.
@@
Experimental setup   We fine-tune two representative open-source LVLMs, InternVL3-8B [68] and Qwen2.5-VL-7B [69]. Training details are reported in Supp. D.1. Our evaluation benchmark consists of two parts: 1) public benchmarks including InfographicVQA [42], which focuses on general infographics with only a subset being infographic charts, and ChartQAPro [41], which covers various chart types; and 2) the aforementioned independent evaluation set of 2,176 charts with 4,975 question-answer pairs specifically targeting infographic charts. For the evaluation metrics, we follow previous work on chart question answering [41], using relaxed accuracy with a 5% margin for numerical answers, ANLS for textual answers, and exact matching for multiple-choice questions.
Results and analysis   Tables 1 and 2 show the evaluation results on the public benchmarks and our evaluation set. After fine-tuning with ChartGalaxy, both models demonstrate improved performance gains across all question types. On the public benchmarks (Table 1), InternVL3 improves performance by 3.80% on InfographicVQA and 5.98% on ChartQAPro, while Qwen2.5-VL shows a 4.44% gain on InfographicVQA and a 3.59% improvement on ChartQAPro. On our evaluation set (Table 2), both models show consistent improvements across all question types, with overall gains of +26.87% for InternVL3 and +23.85% for Qwen2.5-VL. The most notable improvements are observed in the visual understanding questions, with increases of up to +60.49% for style detection and +40.78% for visual encoding analysis. These results indicate that existing pre-training routines may underrepresent questions involving chart visual styles and data encoding, an area our dataset helps to supplement. Performance also improves across the text-based and visual-element-based reasoning questions. Qwen2.5-VL performs well on text-based reasoning, while InternVL3 shows relatively stronger gains on visual-element-based reasoning.
@@ 4.2 Benchmarking Infographic Chart Code Generation
Benchmark construction   The benchmark is designed to evaluate the Direct Mimic task [40], where an LVLM is prompted to generate the D3.js code for a given infographic chart image. Due to variation in coding styles and implementation strategies [70], directly comparing code quality is challenging. Therefore, we evaluate the rendered output instead of the code itself. Specifically, we render the output as both an SVG and a PNG: the SVG enables fine-grained analysis, as it contains precise information about visual and textual elements (e.g., positions, colors), while the PNG supports direct visual comparison. To support this task, we sample 500 infographic charts from ChartGalaxy, covering diverse chart types, variations, and layouts. Each chart is paired with a ground-truth triplet: a PNG image, an SVG, and the corresponding tabular data. Benchmark details are provided in Supp. D.2.
@@
Following previous benchmarks [70, 40], we measure the similarity between the ground-truth chart and the one rendered by the generated code at two levels: a high-level score (overall visual similarity judged by GPT-4o with the PNG images) and a low-level score (average similarity across fine-grained SVG elements). To compute the low-level score, we parse the SVG elements from the rendered chart and the ground-truth one and match them based on attributes such as tag types and positions. This matching is formulated as a linear assignment problem and solved using the Jonker-Volgenant algorithm [71, 70, 72]. Based on the matching results, we compute a low-level score by averaging six metrics: area, text, image, color, position, and size. The area metric captures the ratio of matched element area to the total element area. The text and image metrics assess the similarity of generated text and image elements, respectively. The color, position, and size metrics evaluate visual consistency in these attributes among matched elements. Details of the evaluation metrics are provided in Supp. D.2. As in [40], we also calculate the overall score as the average of the high-level and low-level scores, ranging from 0 to 100. Notably, if the code fails to render the chart, both scores are set to 0.
@@
Experimental setup   We benchmark 17 widely used LVLMs, including 12 proprietary ones and 5 open-source ones, as shown in Table 3. Model configurations and detailed prompts are provided in Supp. D.2 and Supp. E.2, respectively.
@@
Results and analysis   Table 3 presents the results of the 17 LVLMs on our benchmark. We highlight key findings below. First, among proprietary models, Gemini-2.5-Pro achieves the highest overall score of 85.21. Among open-source models, Llama-4-Maverick-17B performs the best with a score of 61.29, outperforming the proprietary GPT-4.1-nano (60.06). However, it still lags behind top-performing proprietary models. Within the GPT-4.1 series, GPT-4.1 and GPT-4.1-mini achieve nearly identical overall scores (80.00 vs. 79.69). However, GPT-4.1 consistently outperforms GPT-4.1-mini across all individual low-level metrics, except the size metric, where it scores notably lower (55.61 vs. 62.85). This drop suggests that GPT-4.1 may struggle to preserve element size accurately, highlighting the need for further investigation. Additional results and analysis are provided in Supp. D.2.
@@
Gemini-2.5-Pro [73] 100.00 90.72 95.69 86.37 87.67 89.23 69.05 86.45 83.97 85.21
GPT-4.1 [74] 100.00 90.58 91.58 86.53 87.52 87.13 55.61 83.16 76.84 80.00
Claude-3.7-Sonnet [75] 100.00 88.96 92.39 77.90 84.78 87.57 67.29 83.15 76.66 79.91
GPT-4.1-mini [74] 99.60 88.21 88.31 79.32 86.43 85.61 62.85 81.79 77.59 79.69
OpenAI-o4-mini [76] 98.80 83.13 79.26 67.53 83.93 84.95 64.07 77.14 74.79 75.97
Gemini-2.5-Flash [77] 96.40 84.52 87.02 73.21 77.81 83.01 62.29 77.98 73.12 75.55
OpenAI-o1 [78] 97.20 85.01 78.07 80.28 81.70 82.35 60.76 78.03 71.35 74.69
OpenAI-o3 [76] 92.40 83.84 82.43 72.15 78.79 81.63 63.44 77.05 71.38 74.22
GPT-4o [79] 99.00 74.86 63.54 34.12 76.60 80.12 56.27 64.25 67.10 65.67
GPT-4.1-nano [74] 100.00 74.97 59.94 36.06 69.26 75.10 47.69 60.50 59.62 60.06
Doubao-1.5-Vision-Pro [80] 97.20 59.05 48.08 36.63 60.09 66.02 40.00 51.65 42.58 47.11
Moonshot-v1-Vision [81] 95.20 60.86 45.54 33.68 60.86 63.33 39.81 50.68 38.11 44.39
@@
Llama-4-Maverick-17B [82] 99.60 76.59 56.37 59.24 69.59 75.39 49.87 64.51 58.06 61.29
Qwen2.5-VL-72B [69] 92.60 73.50 63.22 49.33 65.06 73.14 47.53 61.96 52.21 57.09
InternVL3-78B [68] 95.60 73.36 47.98 62.19 65.76 70.35 43.74 60.56 49.58 55.07
Llama-4-Scout-17B [82] 98.20 71.08 51.81 45.27 63.48 69.95 42.28 57.31 46.50 51.91
Qwen2.5-VL-32B [69] 81.60 60.37 49.37 28.19 54.37 61.37 37.60 48.55 44.42 46.48
@@ 4.3 Example-based Infographic Chart Generation
User study setup   We conducted a user study with 16 experts in design or visualization to evaluate the quality of the generated infographic charts using our method and GPT-Image-1 [83], a state-of-the-art image generation model. The evaluation focused on three key metrics: fidelity (how accurately the data is represented), aesthetics (how appealing the infographic chart is), and creativity (how innovative the design is). Experts were asked to rate 30 pairs of infographic charts generated by our method and GPT-Image-1 based on the same tabular data and reference infographic chart, using a five-point Likert scale (1=poor, 5=excellent) for each metric. Fig. 4 shows examples of the generated infographic charts and the reference. The prompt used for GPT-Image-1 is provided in Supp. E.3.
@@
Results and analysis   Table 4 shows that our method significantly outperforms GPT-Image-1 across all three metrics based on the Wilcoxon signed-rank test on the user rating data (p < 0.01): fidelity (average: 4.63 vs. 2.10), aesthetics (4.14 vs. 2.90), and creativity (3.95 vs. 2.65). Particularly, our method achieves high fidelity by accurately representing data through carefully implemented chart variations. Some scores are slightly lower due to expert preferences for alternative chart types in certain cases. In contrast, GPT-Image-1 often exhibits serious fidelity-related issues, such as incorrect labels, disproportionate representations, and mismatched data elements. In terms of aesthetics and creativity, our method benefits from accurately extracting high-quality layout templates from reference infographic charts and supporting a wider variety of chart types beyond basic chart types, such as bar charts and line charts. By comparison, GPT-Image-1 tends to use basic chart types with limited variations, leading to outputs that are visually simple and monotonous. The detailed analysis is provided in Supp. D.3.
@@ 5 Conclusion
We echo the growing interest in multimodal understanding and generation in LVLMs by introducing ChartGalaxy, a million-scale, high-quality dataset of 104,519 real and 1,151,087 synthetic infographic charts. Grounded in real designs, our structured synthesis pipeline enables the scalable creation of diverse infographic charts. By providing aligned data-chart pairs, extracted layout templates, and three representative applications, we aim to advance the development of foundation models capable of interpreting, reasoning, and generating complex infographic charts.
At the same time, we acknowledge the current ChartGalaxy dataset primarily focuses on single-chart infographics, limiting its ability to capture the complexity of multi-chart narratives. As a result, future work should explore generating and analyzing multi-chart infographics, which emphasize storytelling through coordinated visual elements. Additionally, enriching the interplay between text and visuals could further enhance models’ capacity for nuanced multimodal understanding.

### v5

@@ 3.5 Data Statistics
ChartGalaxy contains 1,701,356 programmatically generated infographic charts and 61,833 real ones, covering 75 chart types, 440 chart variations, and 68 layout templates. Each infographic chart in the dataset is associated with tabular data, providing rich supervision for training and evaluation tasks. Among the 75 chart types, the most frequently occurring ones are horizontal bar charts (11.7%), vertical bar charts (4.9%), and scatterplots (3.5%). Each chart type has up to 27 variations, capturing a wide range of styles in element shapes, icon placement, and rendering effects. We observe differences in the frequency of template usage. For example, the template in Fig. 3(f) is the most common in the synthetic charts (6.9%). Differences in template usage are influenced by each template’s structural flexibility, compatibility with chart types, and prevalence in real-world use cases. For more analysis on our dataset, please refer to Appendix C.
@@ 4.1 Instruction Dataset for Infographic Chart Understanding
In this experiment, we construct an instruction dataset with ChartGalaxy to enhance model capabilities on infographic chart understanding. We validate its usefulness by fine-tuning two open-source LVLMs, demonstrating improvements on both public benchmarks and our independent evaluation set.
Dataset construction To improve LVLMs’ data comprehension and visual understanding of infographic charts, we construct an instruction dataset comprising 443,455 question-answer pairs based on 70,248 charts sampled from ChartGalaxy, ensuring balanced coverage of diverse chart types for better fine-tuning performance. The questions are classified into three types: 1) Text-based reasoning. We incorporate well-established question types from prior work, including open-form questions (Masry et al., 2025) and template-based questions (Meng et al., 2024). These questions cover data identification (DI), data comparison (DC), data extraction with condition (DEC), and fact checking (FC). 2) Visual-element-based reasoning. We extend beyond purely text-based reasoning questions by incorporating visual elements from charts, such as icons (e.g., “What was the wine export value of [Uncaptioned image] in 2024?”). These questions require models to associate visual elements with their corresponding data values, thus testing their ability to conduct more complex cross-modal reasoning (Lin et al., 2025). 3) Visual understanding. This type includes style detection (SD), visual encoding analysis (VEA, e.g., “What data dimension is encoded using different colors in this infographic chart?”), and chart classification (CC). The three types of questions evaluate a model’s ability to interpret visual elements and underlying structural representations of the data. Detailed prompts and methods for generating question-answer pairs are provided in Appendix I.1. Moreover, we construct an independent, human-verified evaluation set containing 2,176 synthetic charts with 4,975 question-answer pairs. We focus on synthetic charts here, as they provide bounding-box annotations that enable visual-element-based reasoning questions.
@@
Experimental setup We fine-tune two representative open-source LVLMs, InternVL3-8B (Zhu et al., 2025b) and Qwen2.5-VL-7B (Bai et al., 2025). Training details are reported in Appendix H.1. Our evaluation benchmark consists of two parts: 1) public benchmarks including InfographicVQA (Mathew et al., 2022), which focuses on general infographics with only a subset being infographic charts, and ChartQAPro (Masry et al., 2025), which covers various chart types; and 2) the aforementioned independent evaluation set of 2,176 charts with 4,975 question-answer pairs specifically targeting infographic charts. For the evaluation metrics, we follow previous work on chart question answering (Masry et al., 2025), using relaxed accuracy with a 5% margin for numerical answers, ANLS for textual answers, and exact matching for multiple-choice questions.
Results and analysis Tables 1 and 2 show the evaluation results on the public benchmarks and our evaluation set. After fine-tuning with ChartGalaxy, both models demonstrate improved performance gains across all question types. On the public benchmarks (Table 1), InternVL3 improves performance by 3.80% on InfographicVQA and 5.98% on ChartQAPro, while Qwen2.5-VL shows a 4.44% gain on InfographicVQA and a 3.59% improvement on ChartQAPro. On our evaluation set (Table 2), both models show consistent improvements across all question types, with overall gains of +26.87% for InternVL3 and +23.85% for Qwen2.5-VL. The most notable improvements are observed in the visual understanding questions, with increases of up to +60.49% for style detection and +40.78% for visual encoding analysis. These results indicate that existing pre-training routines may underrepresent questions involving chart visual styles and data encoding, an area our dataset helps to supplement. Performance also improves across the text-based and visual-element-based reasoning questions. Qwen2.5-VL performs well on text-based reasoning, while InternVL3 shows relatively stronger gains on visual-element-based reasoning. Additional results, including ablation studies on real/synthetic charts and evaluation on a smaller model (Qwen2.5-VL-3B), are provided in Appendix H.1.
@@ 4.2 Benchmarking Infographic Chart Code Generation
Benchmark construction The benchmark is designed to evaluate the Direct Mimic task (Yang et al., 2025), where an LVLM is prompted to generate the D3.js code for a given infographic chart image. Due to variation in coding styles and implementation strategies (Si et al., 2025), directly comparing code quality is challenging. Therefore, we evaluate the rendered output instead of the code itself. Specifically, we render the output as both an SVG and a PNG: the SVG enables fine-grained analysis, as it contains precise information about visual and textual elements (e.g., positions, colors), while the PNG supports direct visual comparison. To support this task, we randomly sampled 500 synthetic infographic charts, with explicit coverage of all chart types, variations, and layout templates. Each chart is paired with a ground-truth triplet: a PNG image, an SVG, and the corresponding tabular data. Benchmark details are provided in Appendix H.2.
@@
Following previous benchmarks (Si et al., 2025; Yang et al., 2025), we measure the similarity between the ground-truth chart and the one rendered by the generated code at two levels: a high-level score (overall visual similarity judged by GPT-4o with the PNG images) and a low-level score (average similarity across fine-grained SVG elements). To compute the low-level score, we parse the SVG elements from the rendered chart and the ground-truth one and match them based on attributes such as tag types and positions. This matching is formulated as a linear assignment problem and solved using the Jonker-Volgenant algorithm (Crouse, 2016; Si et al., 2025; Chen et al., 2024). Based on the matching results, we compute a low-level score by averaging six metrics: area, text, image, color, position, and size. The area metric captures the ratio of matched element area to the total element area. The text and image metrics assess the similarity of generated text and image elements, respectively. The color, position, and size metrics evaluate visual consistency in these attributes among matched elements. Details of the evaluation metrics are provided in Appendix H.2. Following Yang et al. (2025), we also calculate the overall score as the average of the high-level and low-level scores, ranging from 0 to 100. Notably, if the code fails to render the chart, both scores are set to 0.
@@
Experimental setup We benchmark 17 widely used LVLMs, including 12 proprietary ones and 5 open-source ones, as shown in Table 3. Model configurations and detailed prompts are provided in Appendix H.2 and I.2, respectively.
@@
Results and analysis Table 3 presents the results of the 17 LVLMs on our benchmark. We highlight key findings below. First, among proprietary models, Gemini-2.5-Pro achieves the highest overall score of 85.21. Among open-source models, Llama-4-Maverick-17B performs the best with a score of 61.29, outperforming the proprietary GPT-4.1-nano (60.06). However, it still lags behind top-performing proprietary models. Within the GPT-4.1 series, GPT-4.1 and GPT-4.1-mini achieve nearly identical overall scores (80.00 vs. 79.69). However, GPT-4.1 consistently outperforms GPT-4.1-mini across all individual low-level metrics, except the size metric, where it scores notably lower (55.61 vs. 62.85). This drop suggests that GPT-4.1 may struggle to preserve element size accurately, highlighting the need for further investigation. Additional results and analysis are provided in Appendix H.2.
@@
Gemini-2.5-Pro 100.00 90.72 95.69 86.37 87.67 89.23 69.05 86.45 83.97 85.21
GPT-4.1 100.00 90.58 91.58 86.53 87.52 87.13 55.61 83.16 76.84 80.00
Claude-3.7-Sonnet 100.00 88.96 92.39 77.90 84.78 87.57 67.29 83.15 76.66 79.91
GPT-4.1-mini 99.60 88.21 88.31 79.32 86.43 85.61 62.85 81.79 77.59 79.69
OpenAI-o4-mini 98.80 83.13 79.26 67.53 83.93 84.95 64.07 77.14 74.79 75.97
Gemini-2.5-Flash 96.40 84.52 87.02 73.21 77.81 83.01 62.29 77.98 73.12 75.55
OpenAI-o1 97.20 85.01 78.07 80.28 81.70 82.35 60.76 78.03 71.35 74.69
OpenAI-o3 92.40 83.84 82.43 72.15 78.79 81.63 63.44 77.05 71.38 74.22
GPT-4o 99.00 74.86 63.54 34.12 76.60 80.12 56.27 64.25 67.10 65.67
GPT-4.1-nano 100.00 74.97 59.94 36.06 69.26 75.10 47.69 60.50 59.62 60.06
Doubao-1.5-Vision-Pro 97.20 59.05 48.08 36.63 60.09 66.02 40.00 51.65 42.58 47.11
Moonshot-v1-Vision 95.20 60.86 45.54 33.68 60.86 63.33 39.81 50.68 38.11 44.39
@@
Llama-4-Maverick-17B 99.60 76.59 56.37 59.24 69.59 75.39 49.87 64.51 58.06 61.29
Qwen2.5-VL-72B 92.60 73.50 63.22 49.33 65.06 73.14 47.53 61.96 52.21 57.09
InternVL3-78B 95.60 73.36 47.98 62.19 65.76 70.35 43.74 60.56 49.58 55.07
Llama-4-Scout-17B 98.20 71.08 51.81 45.27 63.48 69.95 42.28 57.31 46.50 51.91
Qwen2.5-VL-32B 81.60 60.37 49.37 28.19 54.37 61.37 37.60 48.55 44.42 46.48
@@ 4.3 Example-based Infographic Chart Generation
User study setup We conducted a user study with 16 experts in design or visualization to evaluate the quality of the generated infographic charts using our method and GPT-Image-1 , a state-of-the-art image generation model. The evaluation focused on three key metrics: fidelity (how accurately the data is represented), aesthetics (how appealing the infographic chart is), and creativity (how innovative the design is). Experts were asked to rate 30 pairs of infographic charts generated by our method and GPT-Image-1 based on the same tabular data and reference infographic chart, using a five‑point Likert scale (1=poor, 5=excellent) for each metric. Fig. 4 shows examples of the generated infographic charts and the reference. The prompt used for GPT-Image-1 is provided in Appendix I.3.
@@
Results and analysis Table 4 shows that our method significantly outperforms GPT-Image-1 across all three metrics based on the Wilcoxon signed-rank test on the user rating data (p < 0.01): fidelity (average: 4.63 vs. 2.10), aesthetics (4.14 vs. 2.90), and creativity (3.95 vs. 2.65). Particularly, our method achieves high fidelity by accurately representing data through carefully implemented chart variations. Some scores are slightly lower due to expert preferences for alternative chart types in certain cases. In contrast, GPT-Image-1 often exhibits serious fidelity-related issues, such as incorrect labels, disproportionate representations, and mismatched data elements. In terms of aesthetics and creativity, our method benefits from accurately extracting layout templates from reference infographic charts and supporting a wider variety of chart types beyond basic chart types, such as bar charts and line charts. By comparison, GPT-Image-1 tends to use basic chart types with limited variations, leading to outputs that are visually simple and monotonous. The detailed analysis is provided in Appendix H.3.
@@ 5 Conclusion
We echo the growing interest in multimodal understanding and generation in LVLMs by introducing ChartGalaxy, a million-scale, high-quality dataset of 61,833 real and 1,701,356 synthetic infographic charts. Grounded in real designs, our structured synthesis pipeline enables the scalable creation of diverse infographic charts. By providing aligned data-chart pairs, extracted layout templates, and three representative applications, we aim to advance the development of foundation models capable of interpreting, reasoning, and generating complex infographic charts.
At the same time, we acknowledge that the current ChartGalaxy dataset primarily focuses on single-chart infographics, limiting its ability to capture the complexity of multi-chart narratives. As a result, future work should explore the generation and analysis of multi-chart infographics, which emphasize storytelling through coordinated visual elements. Additionally, enriching the interplay between text and visuals could further enhance models’ capacity for nuanced multimodal understanding.
Ethics Statement
We strictly comply with the terms of use of the original sources and standard research ethics in dataset construction and evaluation. 1. Real infographic charts. We only collect infographic charts from sources that explicitly permit research use: i) 18 chart-rich websites (e.g., Visual Capitalist, Statista), each of which provides clear copyright or license terms (see Table A); ii) Google Images and Bing Images, where we apply built-in license filters to retain only charts explicitly marked under Creative Commons licenses. To further reduce copyright risks, we release only the URLs of real charts, following prior work (Masry et al., 2024; Schuhmann et al., 2022). 2. Synthetic infographic charts. The synthetic infographic charts are generated using resources that are permitted for academic use, including tabular data from openly licensed repositories: UN data (Public domain), Our World in Data (CC BY), VizNet (Hu et al., 2019) (CC BY-NC-SA), and Papers with Code (CC BY-SA); as well as images from publicly available resources with explicit reuse permissions: Icon645 (Lu et al., 2021) (CC BY-NC-SA), Flaticon (Royalty-Free License), Iconshock (Royalty-Free License), Heroicons (MIT License), and Google Material Icons (Apache License 2.0). 3. Sensitive content. To mitigate the risk of exposing sensitive and personal content, we ensure that all real infographic charts, tabular data, and image resources are sourced exclusively from reputable public platforms or collected via search engines with built-in safe-search functionality, both of which generally employ safeguards. We further applied Gemini-2.0-Flash filtering to exclude sensitive topics such as religious conflicts and individually identifiable medical records. 4. Human subjects. The user study (Sec. 4.3) was approved by the Institutional Review Board of the first author’s university. Each participant was compensated with 30 USD for their participation. The study did not involve exposure to emotionally charged, political, or misleading content. Participants were shown infographic charts on neutral topics (e.g., bird population growth, energy sources) and asked to evaluate their quality. No sensitive or personally identifiable data was collected during the study. Participants were fully informed of their rights and were free to withdraw at any time.